OpenStudy (anonymous):

A cylindrical can of milk measures 10 cm high and the diameter of the base is 7 cm. What is the surface area of the can.

OpenStudy (anonymous):

If you can't figure it out first, tell me and I'll help you through it. OK.

OpenStudy (anonymous):

Think of the can as a flat shape. It'll help you.

OpenStudy (gudden):

t.sa. = 2 pie r^2 + 2 pie rh since diameter = 7cm .... it means radius = 7/2 = 3.5 and u already have the height.. :D use... the formula....now..

OpenStudy (anonymous):

Is it multiple choice?

OpenStudy (gudden):

@OrangeMaster looks like..... its not :D

OpenStudy (anonymous):

You're not supposed to provide answers before giving them a chance to work through it themselves, lol, so let's not confirm that just yet.

OpenStudy (anonymous):

They're not even here any more. I'll show the work I used to get there.

OpenStudy (gudden):

hmn,.... what a help u did... by just giving the answer...:p

OpenStudy (anonymous):

You can help them when they come back.

OpenStudy (gudden):

actually if u see @OrangeMaster.... u haven't given him a chance even to think over it... and @KinzaN .... i agree with u that he's not supposed to give the answer...

OpenStudy (anonymous):

If the can of milk has a base and a top, then that answer is wrong anyway.

OpenStudy (anonymous):

So it's fine. :)